OpenStudy (anonymous):

if i simplayfie 12/36 will it be 1/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes good job

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

you're welcome!

OpenStudy (cwrw238):

yes because 36 / 12 = 3

OpenStudy (anonymous):

Yes because they can both be divided by 12. 12/12 = 1 and 36/12 = 3. So it equals 1/3